I'm thinking of a number that, when you multiply it by 9, it is the same as adding 8 to 7 times the number. What is the number?

Answer by ankor@dixie-net.com(22740)

I'm thinking of a number that, when you multiply it by 9, it is the same as adding 8 to 7 times the number.
What is the number?
:
9x = 8 + 7x
9x - 7x = 8
2x = 8
x = 4
:
:
See if that is so:
9(4) = 8 + (7*4)
36 = 8 + 28